### v2

# Multi-Head RAG: Solving Multi-Aspect Problems with LLMs

Abstract
Retrieval Augmented Generation (RAG) enhances the abilities of Large Language Models (LLMs) by enabling the retrieval of documents into the LLM context to provide more accurate and relevant responses. Existing RAG solutions do not focus on queries that may require fetching multiple documents with substantially different contents. Such queries occur frequently, but are challenging because the embeddings of these documents may be distant in the embedding space, making it hard to retrieve them all. This paper introduces Multi-Head RAG (MRAG), a novel scheme designed to address this gap with a simple yet powerful idea: leveraging activations of Transformer’s multi-head attention layer, instead of the decoder layer, as keys for fetching multi-aspect documents. The driving motivation is that different attention heads can learn to capture different data aspects. Harnessing the corresponding activations results in embeddings that represent various facets of data items and queries, improving the retrieval accuracy for complex queries. We provide an evaluation methodology and metrics, multi-aspect datasets that we release online, and real-world use cases to demonstrate MRAG’s effectiveness, showing improvements of up to 20% in relevance over standard RAG baselines. MRAG can be seamlessly integrated with existing RAG frameworks and benchmarking tools like RAGAS as well as different classes of data stores.
1 Introduction
Large Language Models (LLMs) transformed many machine learning tasks using in-context learning abilities. They achieved such accuracy by leveraging an increasing number of parameters, which in recent models have grown to hundreds of billions, making LLM training expensive in terms of both time and resources. It also comes with the danger of leaking confidential data into model weights (Yan et al., 2024; Shi et al., 2024; Patil et al., 2024). Additionally, continuous training through fine-tuning is necessary to keep LLMs up-to-date. Even using the newest data, LLMs display an ongoing problem of hallucinations (Zhang et al., 2023; Xu et al., 2024c; Huang et al., 2023) by providing factually incorrect information. Retrieval Augmented Generation (RAG) was proposed (Lewis et al., 2020; Guu et al., 2020) in order to address these issues as well as others and make LLMs more trustworthy.
The key idea behind RAG is to enhance the generative model’s capabilities by integrating a retrieval system that fetches relevant passages from a large corpus of data. In this setting, when a query is received, the retrieval system first identifies and retrieves pertinent information, which is fed into the generative model’s context for a more accurate and relevant response. Instead of the model storing information within its weights, RAG effectively leverages external knowledge, reducing hallucinations (by grounding the LLM reply in reliable sources), and ensuring that responses contain up-to-date knowledge (e.g., by accessing the Internet), all without requiring expensive training.
More specifically, there are two main stages in a RAG pipeline: data preparation and query execution. During data preparation, one constructs a vector database (DB) populated with embeddings and their corresponding data items such as documents. During query execution, one constructs an embedding of that query and retrieves data items in the store with similar embeddings.
Intense recent research efforts have been put into RAG (Gao et al., 2024; Zhao et al., 2024; Hu & Lu, 2024; Huang & Huang, 2024; Yu et al., 2024; Mialon et al., 2023; Li et al., 2022). On one hand, different RAG designs have been proposed, for example RAPTOR (Sarthi et al., 2024), Self-RAG (Asai et al., 2023), Chain-of-Note (Yu et al., 2023), and many others (Abdallah & Jatowt, 2024; Delile et al., 2024; Edge et al., 2024; Manathunga & Illangasekara, 2023; Zeng et al., 2024; Wewer et al., 2021; Xu et al., 2024b). In general, these schemes focus on making the retrieved data more accurate and relevant to the query. There have also been efforts into benchmarking and datasets for RAG evaluation (Chen et al., 2024b; Xiong et al., 2024; Lyu et al., 2024; Es et al., 2023).
Despite all these advances, we observe that no existing RAG scheme or evaluation methodology explicitly targets an important class of problems that come with a high degree of multi-aspectuality. These are problems that require combining several (potentially many) significantly different aspects in a single query. As a simple illustrative example of such a query, consider the question “What car did Alexander the Great drive?”, and assume that the queried model has not been trained on history. When using RAG, to answer this question accurately, one would retrieve two documents, one describing Alexander the Great and one outlining the history of car manufacturing. However, the embeddings of these two documents could be far away from each other in the embedding space. At the same time, such queries are common in different industry settings, as indicated by extensive discussions with our industry collaborators. Imagine a chemical processing plant experiencing an equipment accident. One could use an LLM to find the accident cause, which might require the retrieval of multiple, potentially confidential documents to provide the necessary context. These documents could be related to different aspects, for example psychological profiles of workers (“Was the accident due to mismanaging a worker?”), equipment purchase records (“Was some equipment part too old?”), maintenance (“Was some equipment part rusty?”), weather (“Was there a particularly strong thunderstorm at the accident time that could have caused dangerous power spikes in the grid?”), or even microclimate (“Was it too humid for an extended period of time in the production hall?”). As we illustrate in Section 4, such problems pose challenges for existing RAG schemes and have been unaddressed by modern RAG benchmarking pipelines.
Figure 1: An overview of the decoder architecture, and a comparison of how standard RAG and Multi-Head RAG embeddings are generated.
In this work, we propose Multi-Head RAG (MRAG): a scheme that addresses the above problem. Common practice in modern RAG designs is the use of embeddings based on last-layer decoder block activations. Our key idea is to use instead the activations of the multi-head attention part of the decoder block as embeddings. The Transformer architecture can be seen as a pipeline with many (e.g., 96 for GPT-3 (Wang et al., 2024b)) blocks, where a single block consists of an attention module and a feed-forward module. Each individual attention module is multi-headed: it consists of multiple parts called heads that learn different sets of weight matrices; see Figure 1 for an overview. It is conjectured that these different heads could capture different aspects of the processed data. We use this as a driving design feature that facilitates capturing the potential multi-aspectuality of the data without increasing space requirements compared to standard RAG, and without any fine-tuning or other modifications to the harnessed model (contribution 1).
Such multi-aspect embeddings are then directly used for both data items and query representation. Considering multi-aspectuality explicitly comes with challenges. For example, it is unclear how to assess whether a RAG solution does indeed harness multiple aspects when fetching documents. For this, we establish an evaluation methodology as well as a full data construction and query processing pipeline that implements the multi-aspect embedding idea (contribution 2). Our datasets facilitate broad evaluation by considering both fully-automatically generated, synthetic data as well as data based on specific industry use cases that show the benefits of MRAG (contribution 3). We ensure the relevance of our RAG datasets in real use cases by working directly with tech leaders (e.g., a generative AI division head) from 3 corporations, all of which actively use RAG in their own LLM infrastructures. Our evaluation illustrates the benefits in the relevance of retrieved documents, for example 20% over a modern RAG baseline for fetching multi-aspect Wikipedia articles, and comparable performance for single-aspect queries (contribution 4). We also show how MRAG and its benchmarking principles can be seamlessly integrated with both existing RAG solutions and benchmarking frameworks such as RAGAS (contribution 5). MRAG’s code is publicly available [^1] https://github.com/spcl/MRAG.
2 The MRAG Formulation & Pipeline
We now present in detail the mathematical underpinning of MRAG and its corresponding pipeline.
Decoder Formulation We first introduce formally the decoder architecture. We omit, for clarity, unnecessary details such as layer normalizations. The input is a text chunk that consists of $n$ tokens. The output of attention head $h$ for the $i$th token $x_{i}$ is defined as follows (Vaswani et al., 2017):
$\text{head}^{h}(\mathbf{x}_{i})=\sum_{j}w_{ij}\mathbf{v}_{j}^{h},\quad\text{where}$ (1)
$\vspace{-0.5em}w_{ij}=\text{softmax}\left(\left(\mathbf{q}_{i}^{h}\right)^{T}\mathbf{k}_{j}^{h}\right),\quad\mathbf{q}_{i}^{h}=\mathbf{W}_{q}^{h}\mathbf{x}_{i},\quad\mathbf{k}_{j}^{h}=\mathbf{W}_{k}^{h}\mathbf{x}_{j},\quad\mathbf{v}_{j}^{h}=\mathbf{W}_{v}^{h}\mathbf{x}_{j}$ (2)
where $\mathbf{W}_{q}^{h},\mathbf{W}_{k}^{h},\mathbf{W}_{v}^{h}$ are, respectively, learnable query, key, and value projections associated with head $h$, and $\mathbf{x}_{j}$ is the vector embedding of the $j$th token $x_{j}$. These outputs get combined to form the output of the $i$th multi-head attention block as follows:
[…]
where matrix $\mathbf{W}_{o}$ is the linear layer that combines the outcomes of all the attention heads. This step is then followed by the Transformer feed-forward layer.
[…]
Multi-Head RAG Formulation The key idea behind MRAG is simple: instead of the single activation vector generated by the last feed-forward decoder layer for the last token, we harness the $H$ separate activation vectors generated by the last attention layer for the last token, before processing it via $\mathbf{W_{o}}$. This can be formulated as a set of embeddings $\mathcal{S}$:
$\mathcal{S}=\{\mathbf{e}_{k}\forall_{k}\}\quad\text{where}\quad\mathbf{e}_{k}=\text{head}^{k}(\mathbf{x}_{n})$ (4)
which is simply the set of all outputs from the attention heads on the last token $\mathbf{x}_{n}$ of the input. As processing with multiple heads does not change the size of the output vector, $\mathcal{S}$ has the same space requirements as standard RAG. However, because we capture the separate embeddings before their mixing with $\mathbf{W}_{o}$, we conjecture that it gives more information about what the different parts of the input attend to, facilitating capturing multi-aspectuality.
Naming We use the terms “single-aspect embedding” and “multi-aspect embedding” to refer to, respectively, a small embedding extracted from a single attention head and a collection of all single-aspect embeddings extracted from an attention layer.
2.1 Overview of the Multi-Head RAG Pipeline
We now describe how the above embedding model fits the RAG pipeline. Figure 2 shows a summary of the design. The MRAG pipeline consists of two main parts, dedicated to data preparation [Uncaptioned image] and query execution [Uncaptioned image]. Both parts heavily use the data store [Uncaptioned image] (vector DB).
2.1.1 Data Preparation
When preparing data [Uncaptioned image], we populate a data store [Uncaptioned image] with multi-aspect MRAG text embeddings [Uncaptioned image] and their corresponding documents [Uncaptioned image] or text chunks [Uncaptioned image] (MRAG is orthogonal to the type of data being embedded, and while we primarily use chunking of documents in order to reflect modern RAG pipelines, one can also embed whole documents or even other types of data). We create the multi-aspect embedding [Uncaptioned image] of each text chunk [Uncaptioned image] using a selected decoder-based embedding model [Uncaptioned image] (this part is detailed in Section 2.2). The user of the pipeline can plug in their model [Uncaptioned image] of choice as well as use their input data. We also offer a dedicated synthetic data generator [Uncaptioned image] that can be used to construct multi-aspect input documents [Uncaptioned image] (we detail this part in Section 3) for evaluation purposes.
MRAG stores data differently than standard RAG, where a single embedding [Uncaptioned image] points to a single text chunk [Uncaptioned image]. For MRAG, each multi-aspect embedding consists of $h$ single-aspect embeddings [Uncaptioned image], each pointing to the original text chunk [Uncaptioned image]. So the data store [Uncaptioned image] contains $h$ embedding spaces, each capturing a different aspect of the text. This crucial feature allows MRAG to compare query [Uncaptioned image] and text chunks [Uncaptioned image] in multiple embedding spaces that capture multiple aspects of the data.
2.1.2 Query Execution
During query execution [Uncaptioned image], we first generate a multi-aspect embedding [Uncaptioned image] of the input query [Uncaptioned image], using the selected embedding model [Uncaptioned image] (details in Section 2.2). Then, we find the nearest multi-aspect embeddings [Uncaptioned image] and their corresponding text chunks [Uncaptioned image] in the data store [Uncaptioned image] using a special multi-aspect retrieval strategy [Uncaptioned image] (detailed in Section 2.3). We ensure that there is no overhead in latency due to multiple aspects because computing these different smaller embeddings is done fully in parallel. Finally, the retrieved data can optionally be assessed [Uncaptioned image] with novel metrics regarding how well it corresponds to the multi-aspect requirements (detailed in Section 3). As with the data preparation [Uncaptioned image] stage, the query execution [Uncaptioned image] stage is flexible, and the user can plug in their models [Uncaptioned image] / [Uncaptioned image] of choice and use their own queries [Uncaptioned image]. We also offer a dedicated synthetic query generator [Uncaptioned image] that can be used to construct multi-aspect input queries [Uncaptioned image] (detailed in Section 3) for evaluation purposes.
Figure 2: Overview of the MRAG pipeline, consisting of two parts: data preparation Refer to caption and query execution Refer to caption. The embedding model Refer to caption and the data store Refer to caption are used by both parts. The data store Refer to caption contains text embeddings Refer to caption linking to text chunks Refer to caption reflecting three different aspects (cyan, magenta, yellow). Blocks marked by a star Refer to caption are a novelty of this work.
2.2 Constructing Multi-Aspect Embeddings [Uncaptioned image]
MRAG can leverage any embedding model with multi-head attention support to construct the multi-aspect embeddings for a given input text. In this work, we consider two embedding models from the MTEB leaderboard (Huggingface, 2024) as potential candidates. Specifically, the SFR-Embedding-Model (Meng et al., 2024) and the e5-mistral-7b-instruct (Wang et al., 2024a), both based on the Mistral 7B architecture with 32 decoder blocks and 32 attention heads per multi-head attention.
While our approach allows for extracting and using the multi-aspect embeddings from any decoder block, and from different layers within a block, we found that multi-aspect embeddings extracted from the last multi-head attention worked best in our experimental setting. We provide further discussion on the carried out experiments in Section 4.
2.3 Retrieval Strategies for Multi-Aspect Data [Uncaptioned image]
A retrieval strategy determines how we select the closest text chunks from the DB given a multi-aspect embedding of the user query. In general, the MRAG retrieval strategy consists of three steps. First, during data preparation, we assign importance scores to all $h$ embedding spaces. Intuitively, these scores capture the fact that different spaces (and the corresponding heads) may be more or less relevant for the used data. Then, during query execution, MRAG starts by applying the traditional RAG retrieval separately for each embedding space. This returns a list of $c$ closest text chunks for each embedding space (a total of $h$ lists). Here, we use a special voting strategy to pick overall top $k$ out of all $hc$ chunks, using the pre-computed importance scores.
Algorithm 1 details the construction of importance scores. It is a heuristic based on extensive empirical evaluation; it gives high-quality results across the tested datasets and tasks. Intuitively, the score $s_{i}$ of a given head $h_{i}$ consists of two parts, $a_{i}$ and $b_{i}$. $a_{i}$ is the average of L2 norms of all embeddings in the vector space $i$; it represents how important a given head is: the larger the norms, the more attention was given to this attention head. $b_{i}$ is the average of cosine distances between all (or a randomly sampled subset, if the user wants to reduce pre-compute time) embeddings in vector space $i$.
Algorithm 1 Importance scores for heads.
  for each head $h_{i}$ do
     $a_{i}\leftarrow 0$; $b_{i}\leftarrow 0$
     $count\_a_{i}\leftarrow 0$; $count\_b_{i}\leftarrow 0$
     for each embedding $e_{ij}$ in $h_{i}$ do
        $a_{i}\leftarrow a_{i}+||e_{ij}||$
        $count\_a_{i}\leftarrow count\_a_{i}+1$
        for each embedding $e_{ih}$ do
           $b_{i}\leftarrow b_{i}+\text{cosine-distance}(e_{ij},e_{ih})$
           $count\_b_{i}\leftarrow count\_b_{i}+1$
        end for
     end for
     $a_{i}\leftarrow a_{i}/count\_a_{i}$; $b_{i}\leftarrow b_{i}/count\_b_{i}$
     $s_{i}\leftarrow a_{i}\cdot b_{i}$
  end for
Intuitively, $b_{i}$ is a proxy for measuring the “spread” of vector space $i$: the larger $b_{i}$, the larger the average angle between different embeddings in this space is. Deriving $s_{i}$ as a product $a_{i}\cdot b_{i}$ ensures that we reward heads with high average attention and high average spread, but simultaneously penalize heads with lower average attention or with low average spread (both $a_{i}$ and $b_{i}$ are appropriately scaled).
The used voting strategy combines the constructed lists of text chunks from individual embedding spaces into a single list of top $k$ chunks. The strategy is very simple (the corresponding Algorithm 2 is in the Appendix). Each text chunk from a list $i$ of the vector space $i$ has a certain position on this list, we denote this position with $p$. We obtain a weight for this chunk as $s_{i}\cdot 2^{-p}$; $s_{i}$ is the previously defined importance score of the space $i$. Multiplying $s_{i}$ with $2^{-p}$ exponentially lowers the significance of less relevant text chunks. Finally, all chunks from all lists are sorted using their weights and the top $k$ chunks form the final list.
2.3.1 Integration with Data Stores & Other Types of Models
MRAG can be seamlessly used with different classes of data stores [Uncaptioned image] and nearest neighbor (NN) search approaches. It can be combined with both the exact and the approximate NN to find the matching (embedding, chunk)-pairs. These two parts of the broader RAG processing pipeline are orthogonal to MRAG. Similarly, MRAG does not depend specifically on the embedding form, as long as it is based on a model that harnesses multi-head attention any such approach that results in a valid embedding can be used. As such, it could also be used with models such as RetroMAE (Xiao et al., 2022) and the classic BGE-embeddings (Xiao et al., 2022; Chen et al., 2024a).
3 Multi-Aspect Datasets, Queries, and Metrics
To assess how well MRAG performs on multi-aspect queries, and to compare it to modern RAG schemes, we need (1) datasets of documents that capture multi-aspectuality, (2) queries to the LLM that touch upon multi-aspectuality and require retrieving different documents from the multi-aspect dataset, and (3) metrics that asses how well a given RAG scheme retrieves such multi-aspect data. We now describe these three elements. In Section 4, we also discuss real-world data and queries used to evaluate MRAG.
3.1 Multi-Aspect Datasets
We first select conceptually different categories of documents for a synthetic dataset. Here, we harness publicly available Wikipedia articles. In the dataset construction pipeline, the user selects a given number of categories (e.g., countries, board games, historical swords, shipwrecks, etc.) and then, for each category, they sample a specified number of documents. The first part of the document (overview) is used as a text chunk to be embedded. We enforce that each overview must have at least 800 characters, matching commonly used chunk sizes in RAG schemes. We also use multi-aspect real-world inspired datasets consisting of NDAs and reports describing industry accidents in chemical processing plants. We ensure the usefulness of these datasets by working directly with tech leaders from 3 corporations that rely on RAG in their in-house LLM-driven report generation and analytics frameworks. Example categories of the legal documents are legal areas (energy law, family law, criminal law, etc.) or document language style (aggressive, mild, neutral, etc.). Examples of accident causes are natural disasters, human mistakes, or lack of proper training. We fully release these datasets to propel RAG research. Details on all three datasets can be found in the Appendix B.2. In our evaluation, we use a total of 13,750 documents.
3.2 Multi-Aspect Query Generation
We also require queries that touch upon a given number of $n$ aspects. For example, a query with 10 aspects must contain a question about 10 different documents from 10 different categories. We create such queries by selecting $n$ categories, sampling a document from each selected category (ensuring there are no duplicates overall), and then generating a story that combines these documents, using an LLM (GPT-3.5 Turbo). We construct 25 queries with 1, 5, 10, 15 and 20 aspects (125 queries in total). An example multi-aspect query sent to the LLM that requires retrieving 10 documents from 10 different categories, is pictured in the top part of Figure 3.
3.3 Metrics
We also design novel metrics to assess how well a given RAG scheme supports multi-aspectuality. For a query $Q$, a used retrieval strategy $S$ (detailed in Section 2.3), and $n$ documents from $n$ categories to retrieve, $Q_{rel}$ denotes the ideal set of documents that should be retrieved for $Q$. Then, $S(Q,n)$ is the set of the actually retrieved documents. We define the Retrieval Success Ratio as $\Xi(Q,n)=\frac{|S(Q,n)\cap Q_{rel}|}{|Q_{rel}|}$, i.e., the ratio of successfully retrieved relevant documents. Moreover, there is a case when a RAG scheme does not retrieve the exact desired document, but it still retrieves successfully some other document from the same category. While less desired, it still increases chances for a more accurate LLM answer following the retrieval. For example, when asking the LLM to determine the cause of an industry accident, fetching the documents in the same category as the accident being queried about, improves the chances for the LLM to give a more relevant answer. To consider such cases, we use another measure, the Category Retrieval Success Ratio or $\Xi_{c}$. It has the same form as $\Xi(Q,n)$ above, with one difference: $S(Q,n)$ is now the set of all the retrieved documents that belong to categories of the ideal desired documents. Finally, to combine these two metrics, we use the Weighted Retrieval Success Ratio $\Xi_{w}$ as $\Xi_{w}=\frac{w\cdot\Xi+\Xi_{c}}{w+1}$. By varying $w$, the user can adjust the importance of exact document matches and category matches. An example of using these metrics to assess how well MRAG and Standard RAG capture multi-aspectuality is pictured in the bottom part of Figure 3.
Figure 3: An example query used to evaluate different RAG strategies. We mention the documents to be fetched in the text and then assess the success ratio of different RAG strategies in finding these documents and their categories. We mark exact document matches Refer to caption , category matches Refer to caption , documents that match a category multiple times Refer to caption , and text segments with no matching document Refer to caption . Finally, we show the weighted success ratio for each strategy, taking a 2:1 weighting (prioritizing the exact article matches).
4 Evaluation
We now illustrate the advantages of MRAG over the state of the art.
Comparison Baselines We consider three main baselines: Standard RAG, Split RAG, and Fusion RAG (Rackauckas, 2024). The first represents a modern RAG pipeline in which each document uses the activations of the last decoder layer as its embedding. The second is a blend between Standard RAG and MRAG. Specifically, it splits the activation of the last decoder layer in the same way as MRAG and applies a voting strategy. The purpose of Split RAG is to show that MRAG’s benefits come from using the multi-head output as embedding and not merely using multiple embedding spaces. Additionally, we consider Fusion RAG (Rackauckas, 2024), an optional mechanism that we harness to further enhance the benefits of MRAG at the cost of additional tokens (detailed in Section 4.3).
We use queries and metrics introduced in Section 3. We use the weighted retrieval success ratio with 2:1 weighting, which considers category matches as relevant but prioritizes the exact document matches. Figure 3 shows an example query and metrics usage. Each query requires retrieving a specific number of documents and the corresponding non-overlapping categories which define the ground truth. We fetch the top $k$ documents from a database, where $k$ is the “total number of documents fetched for a tested RAG scheme” (including potentially mismatches). Among these $k$ documents, we search for matches with the ground truth.
Samples & Summaries Each data point in our plots corresponds to 25 queries. We present the data using standard boxplots to showcase the distribution. Our primary focus is on the average retrieval performance among those 25 queries.
Figure 4: Retrieval success ratio over 25 queries between MRAG and Standard RAG; each query uses 10 different aspects. The top part presents exact document matches, the bottom part presents category only matches (we explain the metrics in Section 3). A histogram is presented for a specific sample to showcase the detailed distribution among the 25 queries (the number of documents fetched for each query is 30).
Figure 5: Relative retrieval improvement of MRAG over Standard RAG across queries with different numbers of aspects and different embedding models (SFR in the left side, e5 in the right side).
4.1 Superior Performance for Multi-Aspect Queries
We start from the query example in Figure 3 and show first the absolute retrieval performance of MRAG over Standard RAG in Figure 4. We fix the number of aspects present in the queries to 10, and vary the total number of retrieved documents from 10 to 30. MRAG consistently outperforms Standard RAG ($>10\%$ increase in the retrieval success ratio on average for exact document matches). Moreover, the retrieval performance increase is even more significant on category matches ($>25\%$ increase in the retrieval success ratio on average). The performance increase is further detailed in the histograms on the right side. Here, for a specific number of documents fetched, MRAG’s histogram indicates a better distribution of retrieval success ratios (across all 25 queries).
Next, Figure 5 shows the relative weighted performance improvement of MRAG with respect to Standard RAG as we vary the number of aspects present in the queries. We show data for two different embedding models (SFR and e5). MRAG consistently outperforms the Standard RAG by 10-20% on average, not only across the number of documents fetched, but also across the number of aspects present in the replies, for both models.
4.2 Comparable Performance for Single-Aspect Queries
We additionally show in Table 1 that MRAG performs on-par with Standard RAG on queries from our multi-aspect dataset where only a single aspect is expected. Hence, our approach does not suffer from significant decrease in performance for single-aspect tasks.
Table 1: Retrieval success ratio (the exact document match) for 25 queries with a single aspect.
[…]
Documents Fetched Multi-Aspect Dataset Legal Dataset Accidents Dataset
SFR e5 SFR SFR
MRAG Standard RAG MRAG Standard RAG MRAG Standard RAG MRAG Standard RAG
1 24/25 25/25 24/25 25/25 24/25 24/25 25/25 25/25
2 25/25 25/25 25/25 25/25 25/25 25/25 25/25 25/25
3 25/25 25/25 25/25 25/25 25/25 25/25 25/25 25/25
4.3 Further Improvements with Additional Tokens
We now show that MRAG can be seamlessly integrated with other RAG approaches: We combine MRAG with Fusion RAG, representing RAG schemes that use an LLM (additional token cost) for more accurate retrieval. Fusion RAG uses an LLM to create a fixed number of questions about the RAG query. Each question is separately applied through an embedding model using Standard RAG. We apply MRAG’s approach to each of these questions and denote the combined scheme as Fusion MRAG. Red plots of Figure 6 show that both Fusion RAG and Fusion MRAG perform better than Standard RAG, on average gaining 10 to 30% in accuracy. Fusion MRAG performs consistently better than pure Fusion RAG, indicating that these optimizations can be combined together. However, both Fusion strategies introduce a greater variance than MRAG and additional costs in terms of compute, latency, and tokens.
4.4 Benefits from Multi-Head Attention Solely
We also compare MRAG to the Split RAG baseline in Figure 6. The blue plots show the relative weighted performance of MRAG and Split RAG over Standard RAG. MRAG performs better than Split RAG, illustrating that its high accuracy is due to the actual multi-head part, and not merely just partitioning the vector and using multiple embedding spaces.
Figure 6: Relative retrieval improvements of MRAG over Standard RAG for the SFR embedding model compared with Split RAG (the blue plots), and the relative retrieval improvements of Fusion MRAG over Standard RAG compared with Fusion RAG (the red plots).
4.5 Real-World Workloads
Figure 7: Average improvement of the retrieval success ratio of MRAG and Split RAG over Standard RAG for two real-world workloads constructing legal documents (left) and discovering causes of industry accidents (right).
To further illustrate advantages of MRAG, we also consider two real-word use cases from in-house industry data analytics projects, namely, the synthesis of legal documents and the analysis of causes of chemical plant accidents. The results are in Figure 7. In the former (the left side), the task is to create a document based on user requirements that may be related to different aspects, for example to the law being considered (e.g., the British or the US one), the subject (e.g., energetic or civil), the style of the document (e.g., aggressive or mild), etc.. This task is executed with RAG that can fetch documents from a database. In the latter (the right side), the task is to discover a cause of an accident. Here, one also wants to retrieve documents from a database that should be used in the LLM context to facilitate discovering the cause of the accident. The causes are grouped in categories such as utility impact due to severe weather, lack of preparedness and planning, incorrect installation of equipment, lack of maintenance, etc.. Similarly to the previous analyses, we measure the retrieval success ratio over corresponding databases. MRAG offers advantages over other schemes.
4.6 Additional Analyses
Figure 8: Evaluation of different voting strategies for MRAG and Split RAG.
We also analyze the impact of using embeddings from different decoder blocks for MRAG (instead of the last one). Here, we consider taking multi-aspect embeddings from three different layers of the embedding model: after the first multi-head attention block, after multi-head attention block 16 (in the middle of the decoder architecture), and the final multi-head attention. We discover that the last multi-head attention performs the best when compared with the Standard RAG.
We also illustrate selected representative data from a long investigation into two additional voting strategies for MRAG. We compare MRAG (1) where only the exponential lowering of significance of selected chunks is applied ($w_{i,p}=2^{-p}$), and MRAG (2) which assigns the weight for each text chunk based on the distance between the particular text chunk ($d_{i,p}$) and the query ($q$) ($w_{i}=\frac{1}{distance(d_{i,p},q)})$. Figure 8 shows that these voting strategies perform worse on average than our selected strategy for MRAG, justifying its design and selection (described in Section 2.3). We also consider two voting strategies for Split RAG, to further deepen the empirical evaluation. Split (1) only uses the exponential lowering of significance ($w_{i,p}=2^{-p}$) and Split (2) which uses the same strategy as MRAG ($w_{i,p}=s_{i}\cdot 2^{-p}$). Figure 8 (on the right) shows that these voting strategies are on-par with each other while being worse than MRAG, further showcasing the advantages of MRAG.
The complexity of the importance score calculation (Algorithm 1) is $\mathcal{O}(n^{2})$ where $n$ is the number of the embedded documents; it is dominated by calculating the pair-wise cosine similarity and the calculation of the norm. Please note that this step needs to be done only once for each dataset and is not a bottleneck.
Finally, in addition to the performance evaluation, we also investigated the attention heads of the SFR-Embedding-Mistral model as well as LLaMA-2 7B model (model not fine-tuned for text-embedding tasks). This analysis is presented in Appendix C.
5 Related Work
Our work touches on many areas which we now briefly discuss.
Many RAG schemes appeared recently (Gao et al., 2024), using the output of the last decoder layer for embedding generation. In contrast, MRAG leverages different embedding spaces of attention heads to focus on different aspects of documents and queries. As such, it can be combined with other schemes to further improve RAG pipelines.
Retrieval is sometimes enhanced by a cross-encoder reranking phase (Rosa et al., 2022; Nogueira & Cho, 2020; Nogueira et al., 2020; Li et al., 2021; Gao et al., 2021; MacAvaney et al., 2019). In such solutions, typically after retrieving a set of relevant chunks, they are re-ranked using specialized models. In this work, we focus solely on the first retrieval phase, so MRAG can be seamlessly used in conjunction with such cross-encoders.
Structure-enhanced RAG schemes employ different strategies for structuring text to improve retrieval quality. A common idea is to construct a Knowledge Graph from text, which enables retrieval amongst entities and relationships (Jiang et al., 2024; Delile et al., 2024; Hussien et al., 2024; Bui et al., 2024; Xu et al., 2024a). RAPTOR (Sarthi et al., 2024) generates multi-level summaries for clusters of related chunks, building a tree of summaries with increasing levels of abstraction to better capture the meaning of the text. Graph RAG (Edge et al., 2024) creates a Knowledge Graph, and summarizes communities in the graph, which provide data at the different levels of abstraction. All these systems try to improve RAG quality by utilizing additional structures that describe entity relationships or the inner organization of text. Usually, they need a sophisticated preprocessing phase to prepare such structures. MRAG achieves the improvement solely based on the embedding model and has no additional storage requirements, and can be combined with any of these schemes.
6 Conclusion
Retrieval Augmented Generation (RAG) is pivotal for democratizing access to accurate and relevant outputs from large language models (LLMs). Enhancing the precision and relevance of these outputs is a critical goal, especially given the challenges posed by queries requiring the retrieval of multiple documents with significantly different contents. These complex queries are common across various domains, but existing RAG solutions struggle because the embeddings of the necessary documents can be far apart in the embedding space, complicating their retrieval.
To address this gap, we introduced Multi-Head RAG (MRAG), a novel scheme that leverages the activations from the multi-head attention layer of decoder models instead of the traditional feed-forward layer. This approach is grounded in the insight that different attention heads can capture distinct aspects of the data. By using these diverse activations, MRAG creates embeddings that better represent the multifaceted nature of data items and queries, thus enhancing the retrieval accuracy for complex, multi-aspect queries. The simplicity and versatility of this idea allow it to be seamlessly integrated into any modern RAG pipeline or data analytics framework.
Our comprehensive evaluation methodology, including specific metrics, synthetic datasets, and real-world use cases, demonstrates MRAG’s effectiveness. The results indicate a significant improvement in the relevance of retrieved documents, with up to 20% better performance compared to modern RAG baselines. This validates MRAG’s potential to handle the intricacies of multi-aspect queries effectively.
Moreover, MRAG proves to be both cost-effective and energy-efficient. It does not require additional LLM queries, multiple model instances, increased storage, or multiple inference passes over the embedding model. This efficiency, combined with the enhanced retrieval accuracy, positions MRAG as a valuable advancement in the field of LLMs and RAG systems. By addressing the challenges of multi-aspectuality in queries, MRAG paves the way for more reliable and accurate LLM applications across diverse industries.
Acknowledgements
We thank Hussein Harake, Colin McMurtrie, Mark Klein, Angelo Mangili, and the whole CSCS team granting access to the Ault and Daint machines, and for their excellent technical support. We thank Timo Schneider for help with infrastructure at SPCL. This project received funding from the European Research Council (Project PSAP, No. 101002047), and the European High-Performance Computing Joint Undertaking (JU) under grant agreement No. 955513 (MAELSTROM). This project was supported by the ETH Future Computing Laboratory (EFCL), financed by a donation from Huawei Technologies. This project received funding from the European Union’s HE research and innovation programme under the grant agreement No. 101070141 (Project GLACIATION). We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017103.
Appendix
Appendix A Model Design: Additional Details
A.1 Retrieval Strategies for Multi-Aspect Data [Uncaptioned image]
Algorithm 2 Voting strategy.
  $l\leftarrow[]$
  for each head $h_{i}$ and its score $s_{i}$ do
     find best matching $k$ text chunks
     for each chunk $d_{i,p}$ with index $p$ in top $k$ do
        $w_{i,p}\leftarrow s_{i}\cdot 2^{-p}$
        add tuple $(d_{i,p},w_{i,p})$ to $l$
     end for
  end for
  sort $l$ using weights $w_{i,p}$; return top $k$ elems
Appendix B Evaluation Methodology: Additional Details
B.1 Compute Resources
Our experiments were executed with compute nodes containing 4x NVIDIA GH200 and a total memory of 800 GB. In general one GPU with at least 40GB of memory should suffice. We used at most 50GB of storage and the OpenAI API as an external resource. The full experiments took at most three hours of GPU time and the cost for the OpenAI API were at most \$15. We carried out additional experiments, which amounted to around 20 hours of GPU time and cost of \$25 for the OpenAI API. Additional evaluation was executed with a mix of compute resources including NVIDIA A100 and V100 GPUs.
B.2 Dataset Details
Table 2: Overview of the structure and the number of documents in the respective datasets.
dataset #categories #topics #documents total #documents
Wikipedia 25 50 documents per category 1250
Legal Documents 25 25 per category 10 per topic 6250
Accident Reports 25 25 per category 10 per topic 6250
B.3 Prompt Template for the Synthetic Dataset Generation
Table 3: Prompt template for query generation.
Please create a story about the attached <number of articles> articles on the topics <list of titles>.
It is very important that each of the attached articles is relevant to the story, in a way that references the content of the article, not just its title. But please also mention each title at least once. Please make sure that all of the attached articles are relevant to your story, and that each article is referenced in at least two sentences! They do not necessarily have to be referenced in the same order, but make sure no article is forgotten.
Important: Output only the story, no additional text. And do not use bullet points, or paragraphs.
Articles:
Article <title>:
<body>
Again, make sure that you reference all the following topics in your story: <list of titles>
Appendix C Attention Head Analysis
We investigated the attention heads of two models in detail: LLaMA-2 7B and SFR-Embedding-Mistral. We selected these two models for a detailed investigation because the former represents models that are not fine-tuned for text embeddings, while the latter is specifically the text embedding model that we used for our experiments. For each model, we looked specifically at the attention scores within each attention head, i.e., how much attention each head pays to each input token during the inference. Knowing the semantics of the input tokens enables then deriving certain conclusions about multi-aspectuality and attention heads.
We plot selected results in Figure 9. Each heatmap shows the dot-product between key- and value-projections inside a given specified attention head, where line $i$ of a heatmap for attention head h indicates the dot-products between the query-projection of token $i$ and the key-projections of all previous tokens $j<i$ (both models use causal attention).
(a) Head 4 LLaMA-2
(b) Head 8 SFR
(c) Head 22 LLaMA-2
(d) Head 21 SFR
Figure 9: Heatmap plots for selected attention heads of the LLaMA-2 7B and SFR-Embedding-Mistral models.
For both models, we found out that the attention patterns vary significantly between the different attention heads. Still, we encountered two distinct patterns. First, the diagonal lines in Figures 9(a) and 9(b) indicate that, when processing a certain input token x, elevated attention is paid to some tokens that came a constant numbers of steps before x. We postulate that this pattern is likely beneficial to understanding the overall rhythm of a natural language, allowing the model to better identify which words are semantically connected, and which parts of the input text refer to each other. Second, horizontal and vertical lines in Figures 9(c) and 9(d) show that these heads learned to pay attention to specific tokens, regardless of how far apart they are within the input sequence. An intuitive justification for such patterns is the focus on certain semantic aspects of the input sequence.
(a) Attention Head 0 LLaMA-2 7B
(b) Attention Head 8 SFR
(c) Attention Head 14 LLaMA-2 7B
(d) Attention Head 21 SFR
Figure 10: Attention scores for selected attention heads of the LLaMA-2 7B and SFR-Embedding-Mistral models.
We also detail attention scores (after applying softmax) of selected heads in Figures 10 and 11, when the model is processing the last token of its input. We see that some tokens gather a lot of attention from most heads, yet there is always a plethora of passages which are attended differently by any two attention heads. An interesting pattern we encountered was that for the SFR-Embedding-Mistral model (see Figure 11), all heads’ attention spiked significantly on the first line-break in the input sequence - either positively or negatively. We conjecture that this is a consequence of how the embedding model was fine-tuned and its intended usage pattern: embedding queries are usually prepended with a retrieval instruction, which is terminated by a line-break. The model likely learnt to summarise the necessary information about this instruction inside the terminating line-break.
Figure 11: Attention scores for all attention heads of the SFR-Embedding-Mistral model.